OpenStudy (zohaibtarar):

An external circuit with a total series resistance of 50 ohm is connected to a 1.5 V AA battery which has an internal resistance of 1 ohm. What is the battery’s terminal potential difference?

hartnn (hartnn):

do you know which equation to use?

OpenStudy (zohaibtarar):

not sure but i think its I = EMF / (R + r)

hartnn (hartnn):

yes thats correct

hartnn (hartnn):

emf = 1.5 R = 50 r = 1 just plug these in

OpenStudy (zohaibtarar):

29.412 mA thats it? or is there any other step left?

hartnn (hartnn):

Thats just the current flowing through the circuit, The voltage drop (PD) across the battery, can now be found using ohm's law. V_battery = V- I * r_battery r_battery =1 V= 1.5 I= 0.029

OpenStudy (zohaibtarar):

so final answer is 1.49 volt? correct?

hartnn (hartnn):

take 0.029 approximately as 0.03 1.5 - 0.03*1 = 1.5 - 0.03 = 1.47V thats it! :)

OpenStudy (zohaibtarar):

and is there any need of diagram in it? if so can you roughly make one to give me an idea? Thank you so much

hartnn (hartnn):